OpenStudy (anonymous):

solve the proportion for u u/10 = 3/11

OpenStudy (anonymous):

@KendrickLamar2014

OpenStudy (kendricklamar2014):

First, we Simplify both sides of the Equation to get:\[\left(\begin{matrix}1 \\ 10\end{matrix}\right) u = \left(\begin{matrix}3 \\ 11\end{matrix}\right)\]

OpenStudy (kendricklamar2014):

Second, we divide both sides by 1/10:\[\left(\begin{matrix}1/10u \\ 1/10\end{matrix}\right) = \left(\begin{matrix}3/11 \\ 1/10\end{matrix}\right)\]

OpenStudy (kendricklamar2014):

Which equals:\[u = \left(\begin{matrix}30 \\ 11\end{matrix}\right)\]

OpenStudy (kendricklamar2014):

and that's your answer ^

OpenStudy (anonymous):

that's what I did and I though I did it wrong because it won't let me type in fractions, it says to round to the nearest tenth

OpenStudy (kendricklamar2014):

\[\left(\begin{matrix}30 \\ 11\end{matrix}\right) = 2.727...\]

OpenStudy (anonymous):

that's what I thought =/ she told me I worked the problem out wrong =p thank you!!!!